OpenStudy (anonymous):

Find the probability of drawing a jack of hearts from a fair deck of 52 cards with a single draw. Answer 1/4 4/13 1/52 1/13

OpenStudy (anonymous):

Wouldn't it be 4/13

OpenStudy (anonymous):

because It can't be 1/52

OpenStudy (kira_yamato):

there is 1 J <3 So P(J <3) = 1/52 If you like you can view it this way P(J) = 4/52 P(heart) = 1/4 P(J Heart) = 1/4 * 4/52 = 1/52

OpenStudy (anonymous):

1/52? this is mindblowing but thanks!

OpenStudy (anonymous):

well how many jack of hearts do you have in a deck of cards ?

OpenStudy (kira_yamato):

There is only 1 J hearts in the deck of 52 if you see the pic

OpenStudy (anonymous):

Oh I thought it was 4 jacks my fault...